2·)At Nikel, Russia, the annual average concentration of sulfur dioxide is observed to be 50 μg/m, at 15°C and 1 atm. What is this concentration of SO2 in parts per billion?
Answer #1

It is known that

1 m3 = 1000 L.

Concentration of SO2 = 50 μg/m3

Volume of air = 1000 L; temperature of air = 15ºC = (15 + 273) K = 288 K; pressure of air = 1 atm.

Molar mass of SO2 = 64 g/mol.

Determine moles of SO2 corresponding to 50 μg/m3.

Moles SO2 = (volume of air)*(50 μg/m3)

= (1 m3)*(50 μg/m3)

= 50 μg

= [(50 μg)*(1 g)/(1.0*106 μg)*(1 mol)/(64 g)]

= 7.8125*10-7 mole.

Moles air = PV/RT

= (1 atm)*(1000 L)/(0.082 L-atm/mol.K)*(288 K)

= 42.3442 mole.

Mixing ratio = (moles SO2)/(moles air)

= (7.8125*10-7 mole)/(42.3442 mole)

= 1.8450*10-8

≈ 18.45*10-9

This simply means that there are 18.45*10-9 moles SO2 in 1 mole air.

A concentration of 1 mole SO2 in 1.0*109 moles air or more clearly, a concentration of 1.0*10-9 moles SO2 in 1 mole air is defined as a concentration of 1 ppb.

Therefore, concentration of SO2 = 18.45 ppb (ans).